# Robust Inverse Graphics via Probabilistic Inference

### v1

@@ 1 Introduction
We propose robust inverse graphics (RIG), where we view the problem through the lens of probabilistic inference. We rely on a pre-trained scene prior, in our case a prior over neural radiance fields (NeRFs), and a weak prior over corruptions—in the case of 3D scene corruptions, another NeRF with a uniform prior over weights. RIG performs full probabilistic inference over both the scene and corruption NeRFs instead of searching for the most probable solution, such as a point estimate obtained via maximum a posteriori (MAP) inference.
@@
Fine-tuning by minimizing reconstruction error is a standard technique in conditional 3D generation (e.g. Chen et al., 2023). In the presence of priors, fine-tuning corresponds to obtaining a MAP point estimate of the latents (in this case, NeRF and corruption parameters). While MAP may yield good point estimates in a multi-view setting, in a single-image setting it yields a “billboard” solution where the corruption NeRF ends up explaining the scene from the conditioned-on view at the expense of the scene NeRF—the corruption NeRF can become a billboard in front of the camera. Below, we will prove that this problem is intrinsic to MAP inference, and empirically show that doing full probabilistic inference doesn’t suffer from this problem.
@@
    We validate the RIG approach on 3D datasets with a number of prior and NeRF representations, across a number of possible corruptions. We theoretically and empirically show that full probabilistic inference produces better results than point estimates on the monocular depth estimation task, and is competitive with direct depth-regression models.
@@
    To enable RIG on diffusion-based priors, we develop reconstruction-guidance with auxiliary latents (ReGAL) and its importance sampling and sequential Monte Carlo generalizations—a class of general-purpose diffusion conditioning methods that is applicable to latent variable models where a subset of the latents is modeled using a diffusion prior. To the best of our knowledge, we are the first to consider diffusion conditioning in this setting.
@@ 2 Method
Given a single image $y$, we would like to infer the underlying 3D scene representation $x$ over which we have a prior $p(x)$. We assume the scene contains a corruption $c$ with a corresponding prior $p(c)$ and that we have a rendering function $R(x,c,\zeta)$ that produces an image given the scene, the corruption and camera parameters $\zeta$. We form the likelihood $p(y|x,c)$ (we omit $\zeta$ for notational clarity) by treating the rendered image as the mean of an independent Gaussian with a constant-variance observation noise. Our approach, dubbed robust inverse graphics (RIG), performs full posterior inference $p(x,c|y)$ to obtain the scene. Figure 1 illustrates the approach.
@@
Full posterior inference is enough In RIG, we perform full posterior inference to obtain the underlying scene $x,c\sim p(x,c|y)\propto p(x)p(c)p(y|x,c)$ which empirically circumvents the billboard solution (Section 6.1). Intuitively, this can be seen as an instance of the mode not being the same as the typical set. The region around the mode where the corruption completely covers the scene has high density but low volume—there aren’t many corruptions that render exactly to the observed image. On the other hand, the posterior takes into account both density and volume, concentrating on regions with high probability mass—there are many non-billboard corruptions that together with a correct scene render to the observed image, although each such solution may have low density.
@@
We use a mean-field parameterization where $x$ and $c$ are independent $q(x,c)=q(x)q(c)$, with each dimension of $x$ and $c$ being parameterized by a separate Gaussian mean and log standard deviation. We run the optimization multiple times and pick the run with the largest ELBO to avoid getting stuck in local optima.
@@ 3 Diffusion scene priors
Denoising diffusion (Ho et al., 2020) has emerged as a powerful alternative to normalizing flows. While it is possible to directly replace the RealNVP used in ProbNeRF with a diffusion-based prior (Dupont et al., 2022, e.g.), diffusion models allow us to tractably increase the dimensionality of our latent representation. A high dimensional latent space enables high fidelity samples and reconstructions. We build on the Single-Stage Diffusion NeRF (SSDNeRF) framework (Chen et al., 2023) to train the scene prior. SSDNeRF optimizes a set of per-training-example latents $\{x_{n}\}$, also known as GLO latents (Bojanowski et al., 2018), the diffusion prior $p_{\phi}(x)$ parameterized by $\phi$, and the likelihood $p_{\psi}(y|x)$ parameterized by $\psi$. See Appendix D for additional details.
@@
The triplane representation uses a UNet denoiser (Ronneberger et al., 2015) which we found to be insufficiently expressive to model larger scenes. Hence we also implement SRT’s set latent representation which doesn’t suffer from this issue. Since SRT doesn’t learn a prior over $x$, we adopt a transformer-based denoiser that is permutation invariant based on Point-E (Nichol et al., 2022).
@@
where $\lambda_{\text{rec}},\lambda_{\text{diff}}>0$ weighting factors. The negative loss can be interpreted as an ELBO ($-\mathcal{L}(\{x_{n}\},\phi,\psi)\leq\mathbb{E}_{p(y)}\left[\log p_{\theta,\psi}(y)\right]$ where the guide distribution is a delta mass on $x_{n}$ if $y=y_{n}$, $q_{\{x_{n}\}}(x|y_{n})=\delta_{x_{n}}(x)$, making the guide’s entropy term in ELBO vanish; the guide is well-defined during training. Once trained, we can discard the latents and use $p_{\phi}(x)p_{\psi}(y|x)$ as the scene prior. See Appendix E for additional details.
@@ 4 Diffusion conditioning with auxiliary latents
In some experiments, we find that ReGAL is not enough as it samples in an “open-loop”, without re-checking the samples against the posterior of interest. To address this problem, we develop an importance sampling (IS) and a sequential Monte Carlo (SMC) generalization of ReGAL, dubbed ReGAL-IS and ReGAL-SMC. Both approaches use importance weights to check the quality of proposed samples against the posterior of interest. Another advantage of both ReGAL-IS and ReGAL-SMC is that they come with convergence guarantees. Both output a weighted set of $K$ samples $\{(x_{k},c_{k}),w_{k}\}_{k=1}^{K}$ which can be used to estimate the posterior expectation of any test function $f$, $\mathbb{E}_{p(x,c|y)}[f(x,c)]$ as $\sum_{k=1}^{K}w_{k}f(x_{k},c_{k})/\sum_{k=1}^{K}w_{k}$. This estimator is consistent: it converges to the posterior expectation with increasing $K$ (Owen, 2013). We can also resample from the weighted empirical distribution $\sum_{k=1}^{K}w_{k}\delta_{(x_{k},c_{k})}(x,c)/\sum_{k=1}^{K}w_{k}$ to get an unweighted sample.
ReGAL with importance sampling. ReGAL-IS (fully described in Appendix B and Algorithm 2) runs ReGAL (Algorithm 1) $K$ independent times and assigns an importance weight $w_{k}$ to each sample $(x_{k},c_{k})$. An importance weight is a ratio between the unnormalized target distribution—in this case formed by multiplying the diffusion prior over all the noised variables times the likelihood $p(z)p(x|z)p(c|x)p(y|x,c)$—and the ReGAL proposal. Unlike in ReGAL, ReGAL-IS’s importance weights are a mechanism for indicating the quality of each sample. Nevertheless, ReGAL-IS is a kind of guess-and-check procedure since it computes the weights for the full ReGAL trace. This makes it prone to bad inference in high dimensions.
@@
ReGAL with sequential Monte Carlo ReGAL-SMC (fully described in Appendix C and Algorithm 3) addresses this problem by decomposing the unnormalized target distribution of ReGAL-IS into a sequence of intermediate target distributions and runs a propose-weigh-resample procedure at each step in order to approximate these target distributions. The resampling step is key to re-allocating computation to traces which are promising as we sample. It is key to choose a sequence of intermediate target distributions so as to approximate a sequence of smoothing posteriors $p(x_{t},c_{t}|y)$. We form this sequence of targets by multiplying the diffusion prior from pure noise up to $t$ and the lookahead likelihood $\tilde{p}(y|z_{t},c_{t}):=p(y|x=x_{\phi}(z_{t};t),c=c_{t})$, similar to reconstruction-guidance conditioning in (7). We use the ReGAL procedure as the proposal distribution since it naturally factorizes into per-step proposals.
@@ 5 Related work
Analysis by Synthesis Using generative models in visual perception is rare in modern systems in part due to performance issues. A number of systems combine NeRFs with structured generative models with nuisance variables (Yuan et al., 2021; Verbin et al., 2021; Park et al., 2023; Ost et al., 2021). We follow this line of work and extend it with priors over NeRFs and probabilistic inference over the joint distribution of scenes and auxiliary parameters.
Diffusion conditioning To the best of our knowledge, we are the first to consider the problem of diffusion conditioning with auxiliary latents: conditioning models where a diffusion model is used as a prior only over a subset of latent variables. Our proposed methods (ReGAL, ReGAL-IS and ReGAL-SMC) bear similarities to existing diffusion conditioning methods. Unlike classifier-free guidance (Ho & Salimans, 2021) and classifier-guided diffusion (Dhariwal & Nichol, 2021) which require retraining for each new likelihood, we focus on the case when this is undesirable or even impossible. A common approach for addressing this problem is reconstruction-guidance (Ho et al., 2022) which ReGAL builds upon. Wu et al. (2023) propose an SMC-based generalization of reconstruction-guidance and is highly related to ReGAL-SMC. ReGAL-SMC pursues a similar strategy of designing intermediate target distributions that approximate smoothing posteriors and proposal distributions that approximate conditional posteriors. However, in the auxiliary latents case, we must additionally choose the form of the proposal over the auxiliary latent (see Equation 8) as well as the lookahead likelihood conditioned on the diffusion latent and the auxiliary latent. The lookahead likelihood can also be viewed as a twisting function in the SMC literature (Lawson et al., 2022, 2023; Heng et al., 2020; Guarniero et al., 2017; Lioutas et al., 2022) which opens up new ways for learning it.
Figure 6: VSD($\downarrow$) histograms across all conditions. Lower is better.
@@ 6 Experiments
The core benchmark we explore is monocular depth estimation in presence of corruptions. Monocular depth estimation is useful in context of autonomous car navigation (Jing et al., 2022) and robotics (Zhou et al., 2019). In this task, the models are given a single RGB image of a scene and must predict a depth image, representing, for each pixel, the distance of the corresponding object surface element from the camera.
@@
For images with volumetric corruptions, we compute VSD relative to the uncorrupted image (and omit the corruption NeRF when rendering the reconstruction). For RGB image comparisons we use the PSNR metric, defined as $-10\log_{10}\mathrm{MSE}$, where MSE is mean-squared-error across pixels and channels. SeeAppendix F for additional details.
@@
DPT $\mathbf{0.30\pm 0.007}$ $0.63\pm 0.0066$ $0.57\pm 0.0074$ $0.49\pm 0.01$
ProbNeRF MAP $\mathbf{0.30\pm 0.0075}$ $0.44\pm 0.0097$ $0.48\pm 0.0093$ $0.46\pm 0.011$
ProbNeRF VI $0.35\pm 0.009$ $0.32\pm 0.0085$ $0.46\pm 0.0099$ $0.36\pm 0.0091$
SSDNeRF ReGAL $0.31\pm 0.0072$ $\mathbf{0.26\pm 0.0062}$ $\mathbf{0.41\pm 0.0093}$ $\mathbf{0.34\pm 0.0091}$
@@
DPT $0.91\pm 0.0037$ $0.90\pm 0.0035$ $0.90\pm 0.0036$ $\mathbf{0.92\pm 0.0031}$
SSDNeRF ReGAL $\mathbf{0.87\pm 0.0034}$ $\mathbf{0.89\pm 0.0034}$ $\mathbf{0.89\pm 0.0035}$ $0.95\pm 0.0028$
Table 1: Mean VSD values across all conditions. Lower is better. The confidence intervals are 3 $\times$ SEM computed from a single trial via bootstrap.
@@ 6.1 MAP vs full posterior inference
Figure 6 and Table 1 summarize the depth reconstruction results. See Appendix G for color reconstruction metrics. For each algorithm and dataset, we aggregate the metrics across all views and all test scenes. While for every algorithm there is considerable variation in the performance of the algorithms, in aggregate we see that probabilistic inference (ProbNeRF VI and SSDNeRF ReGAL) outperform the point estimates and the regression baseline on the corrupted scenes for ShapeNet. For MSN, our prior does not model the data distribution with enough fidelity to resolve the fine details of objects (see Figure 10), and therefore does not decisively outperform the relatively coarse estimate that DPT provides. For FOV estimation in particular, small errors in the estimated FOV parameter lead to large changes in the predicted depths, which is something DPT does not need to contend with due to the calibration procedure we use to evaluate its predictions. Improving the quality of the prior would be a natural area of future improvement.

### v2

@@ 1 Introduction
We propose robust inverse graphics (RIG), where we view the problem through the lens of probabilistic inference. We rely on a pre-trained scene prior, in our case a prior over neural radiance fields (NeRFs) [^1] We require a 3D representation which supports differentiable rendering and is convenient to place a distribution over. NeRFs were the best at fulfilling these requirements at the time of writing, but other approaches could also work, such as Gaussian splatting., and a weak prior over corruptions—in the case of 3D scene corruptions, another NeRF with a uniform prior over its parameters. RIG performs full probabilistic inference over both the scene and corruption NeRF parameters instead of searching for the most probable solution, such as a single set of NeRF parameters obtained via maximum a posteriori (MAP) inference.
@@
Fine-tuning by minimizing reconstruction error is a standard technique in conditional 3D generation (e.g. Chen et al., 2023). In the presence of priors, fine-tuning corresponds to obtaining a MAP estimate of the NeRF parameters of the scene and the NeRF parameters of the observed corruption. While MAP may yield good point estimates in a multi-view setting, in a single-image setting it yields a “billboard” solution where the corruption NeRF ends up explaining the scene from the conditioned-on view at the expense of the scene NeRF—the corruption NeRF can become a billboard in front of the camera. Below, we prove that this problem is intrinsic to MAP inference, and empirically show that doing full probabilistic inference doesn’t suffer from this problem.
@@
    We validate the RIG approach on 3D datasets with a number of prior and NeRF representations, across a number of possible corruptions. We empirically show that full probabilistic inference produces better results than point estimates on the monocular depth
@@
    To enable RIG on diffusion-based priors, we develop reconstruction-guidance with auxiliary latents (ReGAL) and its importance sampling generalization—a class of general-purpose diffusion conditioning methods that is applicable to latent variable models where a subset of the latents is modeled using a diffusion prior. To the best of our knowledge, we are the first to consider diffusion conditioning in this setting.
@@ 2 Method
Given a single image $y$, we would like to infer the underlying 3D scene representation $x$ over which we have a prior $p(x)$. We assume the scene contains a corruption $c$ with a corresponding prior $p(c)$ and that we have a rendering function $R(x,c,\zeta)$ that produces an image given the scene, the corruption and camera parameters $\zeta$. We form the likelihood $p(y|x,c)$ (we omit $\zeta$ for notational clarity) by treating the rendered image as the mean of per-pixel-and-channel independent Gaussians with a constant-variance observation noise. Our approach, dubbed robust inverse graphics (RIG), performs full posterior inference $p(x,c|y)$ to obtain the scene. Figure 1 illustrates the approach.
@@
Full posterior inference as an effective alternative In RIG, we perform full posterior inference to obtain the underlying scene $x,c\sim p(x,c|y)\propto p(x)p(c)p(y|x,c)$ which empirically circumvents the billboard solution (Section 6.1). Intuitively, this can be seen as an instance of the mode not being the same as the typical set. The region around the mode where the corruption completely covers the scene has high density but low volume—there aren’t many corruptions that render exactly to the observed image. On the other hand, the posterior takes into account both density and volume, concentrating on regions with high probability mass—there are many non-billboard corruptions that together with a correct scene render to the observed image, although each such solution may have low density.
@@
We use a mean-field parameterization where $x$ and $c$ are independent $q(x,c)=q(x)q(c)$, with each dimension of $x$ and $c$ being parameterized by a separate Gaussian mean and log standard deviation. We optimize these parameters using stochastic gradients of the ELBO, estimated via the path derivative estimator (Roeder et al., 2017). We run the optimization multiple times (8 in our experiments) and pick the run with the largest ELBO to avoid getting stuck in local optima (cf. Multi-IVON from Shen et al., 2024).
@@ 3 Diffusion Scene Priors
Denoising diffusion (Ho et al., 2020) has emerged as a powerful alternative to normalizing flows. While it is possible to directly replace the RealNVP used in ProbNeRF with a diffusion-based prior (e.g. Dupont et al., 2022), diffusion models allow us to tractably increase the dimensionality of our latent representation. A high dimensional latent space enables high fidelity samples and reconstructions. We build on the Single-Stage Diffusion NeRF (SSDNeRF) framework (Chen et al., 2023) to train the scene prior. SSDNeRF optimizes a set of per-training-example latents $\{x_{n}\}$, also known as GLO latents (Bojanowski et al., 2018), the diffusion prior $p_{\phi}(x)$ parameterized by $\phi$, and the likelihood $p_{\psi}(y|x)$ parameterized by $\psi$. See Appendix D for additional details.
@@
While we found that the triplane representation—which uses a UNet denoiser (Ronneberger et al., 2015)—was sufficient for the ShapeNet dataset, SRT’s set latents performed much better on the more complex MultiShapeNet dataset. Since SRT doesn’t learn a prior over $x$, we adopt a transformer-based denoiser that is permutation invariant based on Point-E (Nichol et al., 2022). In our experiments, we report performance of the triplane representation for ShapeNet and SRT’s set latents for MultiShapeNet.
@@
where $\lambda_{\text{rec}},\lambda_{\text{diff}}>0$ weighting factors. The negative loss, ($-\mathcal{L}(\{x_{n}\},\phi,\psi)\leq\mathbb{E}_{p(y)}\left[\log p_{\theta,\psi}(y)\right]$, can be interpreted as an ELBO where the guide distribution is a delta mass on $x_{n}$ if $y=y_{n}$, $q_{\{x_{n}\}}(x|y_{n})=\delta_{x_{n}}(x)$, making the guide’s entropy term in ELBO vanish; the guide is well-defined during training. Once trained, we can discard the latents and use $p_{\phi}(x)p_{\psi}(y|x)$ as the scene prior. See Appendix E for additional details.
@@ 4 Diffusion Conditioning with Auxiliary Latents
In some experiments, we find that this basic version of ReGAL is not enough as it samples in an “open-loop”, without re-checking the samples against the posterior of interest. To address this problem, we re-interpret ReGAL as an importance sampling algorithm which produces $K$ samples $\{(x_{k},c_{k}),w_{k}\}_{k=1}^{K}$. An importance weight $w_{k}$ is a ratio between the unnormalized target distribution—in this case formed by multiplying the diffusion prior over all the noised variables times the likelihood $p(z)p(x|z)p(c|x)p(y|x,c)$—and the ReGAL proposal. These importance weights are used to check the quality of proposed samples against the posterior of interest. When $K>1$, ReGAL has convergence guarantees consistent with the classical importance sampling results. The weighted set of $K$ samples can be used to estimate the posterior expectation of any test function $f$, $\mathbb{E}_{p(x,c|y)}[f(x,c)]$ as $\sum_{k=1}^{K}w_{k}f(x_{k},c_{k})/\sum_{k=1}^{K}w_{k}$. This estimator is consistent: it converges to the posterior expectation with increasing $K$ (Owen, 2013). We can also resample from the weighted empirical distribution $\sum_{k=1}^{K}w_{k}\delta_{(x_{k},c_{k})}(x,c)/\sum_{k=1}^{K}w_{k}$ to get an unweighted sample. See Appendix B and Algorithm 2 for an extended description of how to incorporate the $K$ hyperparameter into ReGAL.
@@ 5 Related Work
Analysis by Synthesis Using generative models in visual perception is rare in modern systems in part due to performance issues. A number of systems combine NeRFs with structured generative models with nuisance variables (Yuan et al., 2021; Verbin et al., 2021; Park et al., 2023; Ost et al., 2021). We follow this line of work and extend it with priors over NeRFs and probabilistic inference over the joint distribution of scenes and auxiliary parameters. While it is not our main focus, the posterior samples can be used to quantify scene uncertainty (Guo et al., 2020; Rodriguez-Pardo et al., 2023; Hoffman et al., 2023).
Diffusion conditioning To the best of our knowledge, we are the first to consider the problem of diffusion conditioning with auxiliary latents: conditioning models where a diffusion model is used as a prior only over a subset of latent variables. Our proposed method, ReGAL, bears similarities to existing diffusion conditioning methods. Unlike classifier-free guidance (Ho & Salimans, 2021) and classifier-guided diffusion (Dhariwal & Nichol, 2021) which require retraining for each new likelihood, we focus on the case when this is undesirable or even impossible. A common approach for addressing this problem is reconstruction-guidance (Ho et al., 2022) which ReGAL builds upon. Wu et al. (2023) propose an sequential Monte Carlo (SMC)-based generalization of reconstruction-guidance which comes with convergence guarantees like ReGAL with $K>1$. However, it isn’t applicable to models with auxiliary latents. While it is possible to design an SMC version of ReGAL (Appendix C), it didn’t improve performance metrics in our domain.
Figure 6: VSD($\leftarrow$) histograms across all conditions. Lower is better.
@@ 6 Experiments
The core benchmark we explore is monocular depth estimation in presence of corruptions. Monocular depth estimation is useful in context of autonomous car navigation (Jing et al., 2022) and robotics (Zhou et al., 2019). In this task, the models are given a single RGB image of a scene and must predict a depth image, representing, for each pixel, the distance of the corresponding object surface element from the camera. We also evaluate the quality of RGB reconstructions from multiple test views given one view of a scene with corruptions. The source code for many of these experiments is available at https://github.com/tensorflow/probability/tree/main/discussion/robust_inverse_graphics.
@@
For images with volumetric corruptions, we compute VSD relative to the uncorrupted image (and omit the corruption NeRF when rendering the reconstruction). For RGB image comparisons we use the PSNR metric, defined as $-10\log_{10}\mathrm{MSE}$, where MSE is mean-squared-error across pixels and channels.
@@
DPT $0.30\pm 0.007$ $0.63\pm 0.0065$ $0.58\pm 0.0071$ $0.49\pm 0.01$
ProbNeRF MAP $0.30\pm 0.0034$ $0.43\pm 0.0044$ $0.48\pm 0.0042$ $0.46\pm 0.0047$
ProbNeRF VI $0.30\pm 0.0034$ $0.32\pm 0.0039$ $0.45\pm 0.0043$ $0.36\pm 0.004$
SSDNeRF ReGAL (K=1) $0.32\pm 0.0032$ $0.26\pm 0.0027$ $\mathbf{0.40\pm 0.0043}$ $0.49\pm 0.0052$
SSDNeRF ReGAL (K=4) $0.28\pm 0.0029$ $\mathbf{0.25\pm 0.0027}$ $\mathbf{0.40\pm 0.0042}$ $0.43\pm 0.0049$
SSDNeRF ReGAL (K=8) $\mathbf{0.27\pm 0.0029}$ $\mathbf{0.25\pm 0.0027}$ $0.41\pm 0.0042$ $\mathbf{0.34\pm 0.0041}$
@@
DPT $0.91\pm 0.0037$ $0.90\pm 0.0035$ $0.90\pm 0.0035$ $\mathbf{0.92\pm 0.0031}$
SSDNeRF ReGAL (K=1) $\mathbf{0.87\pm 0.0015}$ $\mathbf{0.88\pm 0.0015}$ $\mathbf{0.88\pm 0.0015}$ $0.95\pm 0.0013$
SSDNeRF ReGAL (K=4) $\mathbf{0.87\pm 0.0015}$ $\mathbf{0.88\pm 0.0015}$ $\mathbf{0.88\pm 0.0015}$ $0.95\pm 0.0013$
SSDNeRF ReGAL (K=8) $0.88\pm 0.0015$ $0.89\pm 0.0014$ $\mathbf{0.88\pm 0.0015}$ $0.95\pm 0.0013$
Table 1: Mean VSD values across all conditions. Lower is better. The confidence intervals are 3 $\times$ SEM computed from a five runs via bootstrap.
@@ 6.1 MAP vs full posterior inference
Figure 6 and Table 1 summarize the depth reconstruction results. For each algorithm and dataset, we aggregate the metrics across all views and all test scenes. While for every algorithm there is considerable variation in the performance of the algorithms, in aggregate we see that probabilistic inference (ProbNeRF VI and SSDNeRF ReGAL) outperform the point estimates and the regression baseline on the corrupted scenes for ShapeNet. For MSN, our prior does not model the data distribution with enough fidelity to resolve the fine details of objects (see Figure 10), and therefore does not decisively outperform the relatively coarse estimate that DPT provides. For FOV estimation in particular, small errors in the estimated FOV parameter lead to large changes in the predicted depths, which is something DPT does not need to contend with due to the calibration procedure we use to evaluate its predictions. Improving the quality of the prior would be a natural area of future improvement.
For ReGAL, we vary the number of importance samples $K$. The decorruption quality is weakly dependent on this for most settings, but is is most noticeably beneficial for the FOV estimation task for ShapeNet.
@@
Figure 7: PSNR($\rightarrow$) histograms across all conditions. Higher is better.
ShapeNet
Model Clean Rain Cloud FOV
ProbNeRF MAP $27.08\pm 0.17$ $18.96\pm 0.3$ $17.68\pm 0.25$ $31.88\pm 0.38$
ProbNeRF VI $27.11\pm 0.2$ $23.22\pm 0.29$ $19.54\pm 0.27$ $32.97\pm 0.47$
SSDNeRF ReGAL (K=1) $30.49\pm 0.16$ $\mathbf{26.58\pm 0.22}$ $19.89\pm 0.31$ $35.08\pm 0.32$
SSDNeRF ReGAL (K=4) $30.80\pm 0.15$ $26.47\pm 0.22$ $\mathbf{19.95\pm 0.29}$ $34.44\pm 0.33$
SSDNeRF ReGAL (K=8) $\mathbf{30.87\pm 0.15}$ $26.32\pm 0.23$ $19.83\pm 0.29$ $\mathbf{35.22\pm 0.27}$
MultiShapeNet
Model Clean Rain Cloud FOV
SSDNeRF ReGAL (K=1) $25.37\pm 0.15$ $\mathbf{19.01\pm 0.38}$ $14.40\pm 0.21$ $23.99\pm 0.31$
SSDNeRF ReGAL (K=4) $25.68\pm 0.13$ $16.74\pm 0.35$ $14.65\pm 0.2$ $25.19\pm 0.26$
SSDNeRF ReGAL (K=8) $\mathbf{25.79\pm 0.13}$ $15.65\pm 0.34$ $\mathbf{14.71\pm 0.2}$ $\mathbf{25.39\pm 0.26}$
Table 2: Mean PSNR values across all conditions. Higher is better. The confidence intervals are 3 $\times$ SEM computed from a five runs via bootstrap. Note that the PSNR values for the FOV dataset are inflated because of the white background which takes up a higher fraction of the image for larger fields-of-view values.